Hypothesis Testing: 1. You are given the following information and then asked to do a hypothesis test with a 95% level of con please help me answer questions 1,2, 3.
Answer #1

(1)

(i)

n = 10000

T = 76

s = \sqrt{10000} = 100

\alpha = 0.05

df = 10000 - 1 = 9999

From Table, critical values of t = \pm 1.96

Confidence Interval:

\bar{x}\pm \frac{t\times s}{\sqrt{n}}=76\pm \frac{1.96\times 100}{\sqrt{10000}}=76\pm 1.96=(74.04,77.96)

So,

Answer is:

(74.04, 77.96)

(ii)

Since 76 is included in the Confidence Interval: (74.04, 77.96), we conclude:

It is likely that the population mean is equal to 76.

(2)

(i)

n = 1500

T = 450

s = \sqrt{10000} = 100

\alpha = 0.05

df = 10000 - 1 = 9999

From Table, critical values of t = \pm 1.962

Confidence Interval:

\bar{x}\pm \frac{t\times s}{\sqrt{n}}=450\pm \frac{1.962\times 100}{\sqrt{1500}}=450\pm 5.065=(444.935,455.065)

So,

Answer is:

(444.935 455.065)

(ii)

Since 440 is not included in the Confidence Interval: (444.935 455.065), we conclude:

It is not likely that the population mean is equal to 440.

(3)

The 2 CIs are different because as Sample Size = n is decreased from n = 10000 to n =1500, the width of the Confidence Interval inceases.